mtzmario8612:

Please help me!!! Multiply (5 × 106)(6 × 102). Express the result in scientific notation.

mtzmario8612:

i meant to put (5 x 10⁶)(6 x 10²)

Extrinix:

So to start off, you need to do the exponents \(\sf{(5\times10^6)(6\times10^2)}\) Figure out your exponents e.g. \(\sf{10^6}\) would be "\(\sf{10 \times 10 \times 10 \times 10 \times 10 \times 10}\)" and \(\sf{10^2}\) would be "\(\sf{10\times10}\)" Once we get those you have: \(\sf{(5\times1,000,000)(6\times100)}\) Now we multiply \(\sf{inside}\) the parenthesis. \(\sf{(5,000,000)(600)}\) We can then remove the parenthesis and multiply then two numbers together. \(\sf{5,000,000\times600}\) Just multiply those together and you have your answer. Best of luck, \(\style{font Family:Arial; text-shadow: 2px 2px 3px white, 0 0 10px white, 0 0 5px white;color:#23B7F3;font-size:100px;padding:50px;width:500px;height:100%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\sf\color {} {{-~Akuma~matata}}}\) "what a wonderful phrase"